### v3

# Fine-Grained Knowledge Structuring and Retrieval for Visual Question Answering

Abstract.
Visual Question Answering (VQA) focuses on providing answers to natural language questions by utilizing information from images. Although cutting-edge multimodal large language models (MLLMs) such as GPT-4o achieve strong performance on VQA tasks, they frequently fall short in accessing domain-specific or the latest knowledge. To mitigate this issue, retrieval-augmented generation (RAG) leveraging external knowledge bases (KBs), referred to as KB-VQA, emerges as a promising approach. Nevertheless, conventional unimodal retrieval techniques, which translate images into textual descriptions, often result in the loss of critical visual details. To address these challenges, this study presents two key innovations. First, we introduce fine-grained knowledge units that consist of multimodal data fragments (e.g., text fragments, entity images, and so on) in a structured manner. Rather than merely refining retrieval mechanisms, we prioritize the systematic organization and management of these knowledge units, ensuring that the structuring process itself enhances retrieval quality. Second, we propose a knowledge unit retrieval-augmented generation framework (KU-RAG) that seamlessly integrates fine-grained retrieval with MLLMs. Our KU-RAG framework not only ensures precise retrieval of relevant knowledge but also enhances reasoning capabilities through a knowledge correction chain. Experimental results demonstrate that our approach consistently outperforms existing KB-VQA methods across four benchmarks, achieving an average improvement of approximately 3% and up to 11% in the best case. The source code is available at: https://github.com/zzhang393/KU-RAG.
1. Introduction
Figure 1. Sample VQA Solution with MLLMs: (a) Direct answer without additional knowledge. (b) Single-modality coarse-grained RAG. (c) Our proposal KU-RAG.
Knowledge-based Visual Question Answering (KB-VQA) extends traditional Visual Question Answering (VQA) by incorporating external knowledge to answer questions where image information alone is insufficient (Marino et al., 2019; Lin et al., 2022; Wen et al., 2024). However, traditional methods often face limitations in their ability to perform complex reasoning over both visual content and external knowledge sources, as they typically rely on predefined retrieval mechanisms or specific training data (Wu and Mooney, 2022; Yang et al., 2023).
VQA with MLLMs. Recently, the emergence of multimodal large language models (MLLMs), such as GPT-4 (Achiam et al., 2023) and LLaVA (Liu et al., 2023a), has introduced new possibilities for VQA. Unlike previous methods, MLLMs serve not only as powerful reasoning engines but also as vast knowledge repositories, with information learned from world knowledge during pretraining (Wang et al., 2024; Liu et al., 2023b). This dual capability enables more nuanced answers. However, the knowledge acquired during training is general and (maybe outdated) world knowledge, limiting the model’s ability to respond to domain-specific and update-to-date queries. As shown in Figure 1(a), when using GPT-4 to ask a question about the bridge in the image, it fails to provide an answer due to a lack of relevant knowledge, and LLaVA even hallucinated and provided a “false” answer.
VQA with RAG and MLLMs. At this point, it becomes necessary to employ KB-VQA methods by retrieving information from a database – a process also known as Retrieval-Augmented Generation (RAG) in the context of LLMs (Fan et al., 2024). This typically involves converting images into captions and then performing passage-level retrieval combined with the query. However, this method struggles to handle fine-grained information for question answering, and during the image-to-text modality conversion process, some visual details are inevitably lost. As shown in Figure 1(b), a unimodal, coarse-grained approach fails to retrieve the relevant knowledge.
Intuitively, in order to accurately find the knowledge corresponding to this bridge, it is necessary to identify the corresponding images through its visual features and then look through the information behind it, as illustrated in Figure 1(c).
Our Proposal: VQA with MLLMs and Fine-Grained Structured Knowledge. Following this approach, we propose a “Knowledge Unit” (KU) component to bridge the query and specific knowledge. Unlike traditional RAG methods that primarily focus on optimizing retrieval strategies, our approach centers on designing a structured knowledge representation that naturally connects queries with relevant knowledge. As shown in Fig. 1 (c), the KU consists of the picture and the description of the Karnin Lift Bridge.
Based on it, we propose a Knowledge Unit Retrieval-Augmented Generation (KU-RAG) method, which is a multimodal, fine-grained, zero-shot retrieval approach covering both data storage and retrieval. As shown in the figure, our method matches the image from the question with images in the database, identifying the relevant knowledge (e.g., “Karnin Lift Bridge”). Instead of merely refining retrieval mechanisms, we emphasize the structured management of knowledge units, where the process of organizing and managing knowledge itself optimizes retrieval quality. Finally, we designed a Knowledge Correction Chain (KCC) to assist in answer generation. The KCC integrates retrieved information into the MLLM’s reasoning process and verifies the accuracy of the knowledge generated by MLLMs.
Contributions. Our key contributions are summarized as follows:
  • We introduce the concept of Knowledge Units (KUs), which structure fine-grained multimodal knowledge for efficient retrieval in database systems. By focusing on the structured management of KUs, we enhance retrieval quality beyond just refining retrieval mechanisms.
  • We propose a knowledge unit retrieval-augmented generation (KU-RAG) method, which retrieves fine-grained knowledge units, employs a knowledge correction chain (KCC) during query inference, and achieves zero-shot for combining retrieved knowledge units with MLLMs.
  • We evaluate our approach on multiple KB-VQA benchmarks, demonstrating its effectiveness in improving knowledge retrieval and reasoning capabilities in VQA.
2. Preliminary
2.1. Task Definition
Visual Question Answering (VQA). Given a question $Q$, which consists of an image $Q_{i}$ and a textual question $Q_{t}$ related to the content of the image, the task of VQA is to generate an answer $A$ based on the information available in the image and the text. In this setup, the system aims to understand both the visual and textual aspects of the input and provide a relevant response.
Knowledge-Based Visual Question Answering (KB-VQA). In KB-VQA, the goal extends beyond the basic VQA task by incorporating external knowledge $K$ stored in knowledge bases (KBs) to answer the question. This external knowledge, which can be categorized as either image knowledge $K_{i}$ or text knowledge $K_{t}$, is retrieved based on the question $Q$ and is used to generate a more informed and accurate answer $A$.
2.2. Knowledge Unit
For KB-VQA, the core challenge lies in accurately locating relevant knowledge. As shown in Figure 2, general coarse-grained retrieval methods, such as image-to-image or caption-based text retrieval, often introduce significant noise or cause loss of visual information, making it difficult to retrieve the correct knowledge. Fine-grained retrieval alleviates some of this noise by focusing on predefined text entities or visual objects, but still fails to preserve complete visual context and cannot effectively integrate textual knowledge at the entity level.
To address this, We introduce a new structure called Knowledge Unit (KU). Each KU serves as a knowledge carrier or object generated in combination with the query, such as entities, events, rules, topics, etc., designed to bridge the gap between the query and the database during the actual question-answering process. Figure 3 illustrates examples of KUs constructed as entities and events.
For a piece of knowledge, the three most important factors are its image, its name, and detailed textual knowledge. Therefore, we designed each KU as a triplet, consisting of Knowledge Image ($K_{i}$), Knowledge Name ($K_{n}$), and Knowledge Text ($K_{t}$):
(1) $KU=\{K_{i},K_{n},K_{t}\}$
In the KB-VQA task, image-image or name-name matching is typically used to determine which piece of knowledge a given image belongs to. Hence, we encapsulate $K_{i}$ and $K_{n}$ into the “Matching End” to link the query and the KU. The purpose of KB-VQA often involves querying the knowledge behind an image, so we refer to $K_{t}$ as the “Detail End”. As shown in the third figure of Fig. 2, we take the Karnin Lift Bridge as an example to construct a KU. During actual operation, the user provides an image $Q_{i}$, which is matched through the Matching End against existing images. The system successfully identifies the KU corresponding to the Karnin Lift Bridge. Subsequently, background knowledge related to this KU is retrieved via its Detail End, and the relevant knowledge fragment is located with the help of the query $Q_{t}$.
Our objective is to manage knowledge effectively through the structured organization and manipulation of KUs, thereby enabling seamless bridging between user queries and knowledge. To achieve this, there are two main challenges:
  • How to classify and extract knowledge from a large amount of unannotated text to construct structured KUs (i.e., the KU construction problem);
  • How to support flexible KU operations such as insertion, deletion, and updating (i.e., the KU operation problem).
We present our solutions to these challenges in Section 3.
Figure 2. Illustration of general coarse-grained retrieval, fine-grained retrieval and knowledge unit retrieval for VQA.
3. Knowledge Unit Management
3.1. Construction
Knowledge Predefinition. Firstly, we should determine how to extract the knowledge unit with the application scenario and consider the query and database. For example, in an object recognition QA system, different entities can serve as knowledge units; in an event query system, different events can serve as knowledge units; in a corporate rules and regulations query system, a rule-based knowledge unit should be constructed. It is important to note that knowledge units do not necessarily need to be atomic or as fine-grained as possible. Its division should be determined based on the granularity of the data in the query and the database. For instance, a general animal knowledge QA system may only require the general species of an animal (e.g., “cat”), whereas a specific species QA system may require the specific species name (e.g., “Persian cat”).
Figure 3. The construction of knowledge units with entity and event. The entity list comes from the work of  Hu et al. (2023b), while the events are sourced from E-VQA (Yang et al., 2023).
Knowledge Segment. Since subsequent steps involve the storage of the knowledge unit, storing textual knowledge $K_{t}$ within the detail end at the document level, which is a coarse-grained storage method, is highly detrimental to knowledge retrieval (Chen et al., 2021). Additionally, during reasoning, this can be limited by the LLM’s maximum token capacity, leading to incomplete information. Therefore, we first need to segment the raw data, breaking it down into finer-grained units.
  • Textual data: Similar to general RAG methods, we next segment all text passage $P=(p_{1},p_{2},\dots,p_{n})$ in a knowledge base to obtain the smallest retrieval units. Each passage $P$ contains $n$ sentences, i.e., $P_{k}=(s_{1},s_{2},\dots,s_{n})$. Considering the importance of knowledge coherence in the KB-VQA task, we adopt a combination of sentence splitting and maximum token limit. In each chunk, as many sentences as possible are retained without exceeding the maximum token limit, and the remaining sentences are assigned to the next chunk. That is, $C=(c_{1},c_{2},\dots c_{i})$ , where $c_{j}=(s_{1},s_{2},\dots s_{j})$, $i$ represents the $i$-th chunk, and $j$ represents the $j$-th sentence in the chunk.
  • Image data: For images pertaining to the same piece of information (such as all images within a news article), we directly extract them, treating each image individually.
Knowledge Assembly.
After segmenting the text and generating chunks, the next step is to assemble these units with the unprocessed image information to form a knowledge unit. In simple terms, we assemble this multimodal knowledge into knowledge units by leveraging the original structural properties of the knowledge and performing an inverted index on the text. To facilitate understanding, we illustrate this entity-type and event-type knowledge unit shown in Figure 3.
Storage
Next, we need to store the knowledge contained within the knowledge units. We encode each chunk into a vector using a text encoder and store them in a Faiss database (Douze et al., 2024), which we denote as $D_{t}$. Considering the need to handle multimodal data in the framework and the possibility of longer text within the chunks, we use Long-CLIP (Zhang et al., 2024) as the vector encoder.
(2) $V_{c_{i}}=\text{Encoder}(c_{i})$
(3) $D_{c}=(V_{c1},V_{c2},\dots,V_{cn})$
For the images in the knowledge base, we also encode each image using Long-CLIP to obtain visual features and store them in the Meta Faiss vector database, denoted as $D_{i}$.
(4) $V_{i_{j}}=\text{Encoder}(i_{j})$
(5) $D_{i}=(V_{i_{1}},V_{i_{2}},\dots,V_{i_{n}})$
3.2. Database-level Operation
Since our approach operates at the database level, compared to traditional methods, we must also consider issues related to data management. Additionally, there is no need to retrain the entire framework after adding, deleting, or updating data or knowledge.
Knowledge Addition. When new knowledge is introduced, we directly chunk the corresponding text. The encoded text vectors are then stored in the existing text vector database $D_{c}$. Similarly, for images within the knowledge base, we encode them and store the resulting vectors in the image vector database $D_{i}$.
Knowledge Unit Operation. After performing operations on the raw data, it is necessary to consider whether corresponding actions need to be taken for the knowledge unit.
1) KU Addition and Update: When new raw data is added, it is essential to assess whether a new KU needs to be introduced. This process primarily involves two steps: first, matching the new knowledge with existing KU. If the similarity of the matching results exceeds the threshold $\alpha$, the index containing the new keywords is added to the corresponding KU through keyword matching. If no matching result exceeds the threshold, a new KU is constructed according to the KU construction rules and the new keywords present in the chunk.
2) KU Deletion: After deleting a chunk from the raw data, it is necessary to check whether the related KU is still valid to reduce storage usage. Specifically, after deleting the chunk indexed as $i$, all KUs containing this index should be checked. If a certain KU has an empty detail end (i.e., no remaining values), that KU can be deleted.
4. Knowledge Unit Retrieval Framework
Figure 4. Overview of KU-RAG Framework
In this section, we will introduce our knowledge unit retrieval framework and detail how to achieve knowledge retrieval through knowledge units and apply it to the KB-VQA task.
As shown in Figure 4, our framework is divided into three modules: Knowledge Unit Construction, Knowledge Retrieval, and MLLM Reasoning & Generation. The knowledge unit construction module mainly transforms raw knowledge into knowledge units and stores them in the database, as illustrated in Section 2.2. The knowledge retrieval module processes the original query, matches it with the corresponding knowledge units, finds the relevant knowledge, and integrates it into a structured, MLLM-readable passage. Finally, combining the original question and the retrieved knowledge, the MLLM Reasoning & Generation module analyzes and generates the answer.
4.1. Query Processing
For the user’s input query $Q$ , it is first necessary to preprocess and rewrite it to reduce interference during the retrieval process. To find the region in the image related to the question, we propose a query-aware instance segmentation method. Specifically, we first use YOLO (Redmon et al., 2016) to perform instance segmentation on the image, obtaining n segmented instance objects $O=(o_{1},o_{2},\dots,o_{n})$. We then encode these instances using Long-CLIP, resulting in corresponding vectors $V_{o}=(v_{o_{1}},v_{o_{2}},\dots,v_{o_{n}})$:
(6) $V_{o_{i}}=\text{Encoder}(o_{i})$
Simultaneously, we encode the textual query $Q_{t}$ into a vector and compute the similarity between it and each vector in $V_{o}$ to find the object related to the query:
(7) $V_{q_{t}}=\text{Encoder}(Q_{t})$
(8) $S_{n}=\text{Sim}(V_{q_{t}},V_{o_{i}})$
Here, $S_{n}=(s_{1},s_{2},\dots,s_{n})$ represents the similarity values corresponding to each O . We select the object o with the highest similarity that exceeds a threshold $\gamma$ for subsequent retrieval, with its vector denoted as $V_{q_{i}}$. Of course, sometimes the query may not only be related to one object but also to areas outside the objects or the entire scene in the image. Therefore, if no object meets the criteria or multiple objects meet the criteria, we will encode the entire image and use this encoding for subsequent retrieval:
(9) $V_{q_{i}}=\text{Encoder}(Q_{i})$
4.2. Knowledge Unit Matching
Next, we use the obtained visual features to match the corresponding knowledge unit. We select the top $k$ knowledge unit items with the highest similarity, denoted as the set $KU^{\prime}$, where each $ku^{\prime}$ contains $j$ indices in its detail end.
(10) $KU=\text{Matching}(V_{q_{i}})$
(11) $KU^{\prime}=\text{top-k}(KU)$
With $KU^{\prime}=(ku_{1}^{\prime},ku_{2}^{\prime},\dots,ku_{k}^{\prime})$. The indices of each $ku_{i}^{\prime}$ are represented as $C_{i}=(c_{i,1},c_{i,2},\dots,c_{i,j})$. Finally, we obtain the combined index set of the knowledge unit:
(12) $C_{ku}=(C_{1},C_{2},\dots,C_{n})$
To integrate KU information into the query while highlighting the importance of certain content words, we rewrite $Q_{t}$ as: "$Q_{t}^{\prime}=Q_{t}\text{ [SEP] }\text{[KU name]}\text{ [SEP] }\text{keywords}$", and encode it as:
(13) $V_{q_{t}}^{\prime}=\text{Encode}(Q_{t}^{\prime})$
Here, “KU name" refers to the name of the matching segment of the retrieved KU, and “keywords" is a list of content words extracted from $Q_{t}$, separated by commas. “[SEP]" is a special token used to separate different parts. Next, we combine the features of $Q_{t}$ and calculate the similarity to obtain the top $k$ chunks related to the query, denoted as $C$’:
(14) $C^{\prime}=\text{top-k}(\text{Sim}(V_{q_{t}},C_{ku}))$
with $C^{\prime}=(c_{1}^{\prime},c_{2}^{\prime},\dots,c_{k}^{\prime})$.
5. MLLM Reasoning and Generation
After retrieving the relevant blocks, the next step is to provide the retrieved information to the MLLM to assist with reasoning and generation. The specific steps are as follows:
Modal Aligning and Fusing. First, based on the retrieval results $C^{\prime}$, we find the corresponding knowledge unit $KU^{\prime}$ for each chunk and combine its matching end information to form an image with the structure ‘[Image][[Name][Chunk Text]]’, where the image corresponding to the $i$-th chunk is denoted as $I^{\prime}_{i}$. Notably, if multiple chunks correspond to the same image, to enhance the connection between knowledge and improve processing efficiency, we merge the texts of these chunks into a single image in the format ‘[Image][[Name][Chunk $Text_{1}$]…[Chunk $Text_{n}$]]’.
Images Stitching. Next, we stitch all the images obtained in the previous step to generate a multimodal passage with both image and text information. This multimodal passage $MP$ is as:
(15) $MP=(I^{\prime}_{1},I^{\prime}_{2},\dots,I^{\prime}_{n})$
Knowledge Correction Chain. At this stage, a key challenge is effectively managing the relationships among the “information in the query”, “the knowledge retrieved”, and “the inherent knowledge of the MLLM”, as well as ensuring a fine-grained correspondence between text and images.
In our experiments, we found that when combining the retrieved knowledge with the question for the MLLM to answer, the model tended to prioritize the retrieved information while neglecting its own knowledge. We also attempted to use guiding prompts, such as “Based on your own knowledge first…” and “Focus on the first image and ignore other image…” to encourage the MLLM to consider its own knowledge before referring to the retrieved information, but the results were unsatisfactory (as demonstrated in Section  6).
To address this issue, we design a Knowledge Correction Chain (KCC) that guides MLLMs in reasoning through multi-turn dialogue and reading comprehension. In detail, we first input the question $Q$ to MLLM to obtain the original answer $A_{0}$:
(16) $A_{0}=\text{MLLM}(Q)$
The purpose of this step is to obtain the pure knowledge of the MLLM regarding the query without being influenced by the retrieved information mentioned above. Finally, we input the passage $MP$ into the MLLM with a knowledge correction aware (KC-aware) prompt and get the final answer $A$:
[KC-aware prompt]: The initial answer has already been provided. The new image information may either be related or unrelated to the previous input. If this new information conflicts with the initial answer, please update the response accordingly. If no changes are needed, simply output the initial answer again.
(17) $A=\text{MLLM}(MP,Prompt,(Q,A_{0}))$
In short, the idea of KCC is to shift the MLLM’s focus from analyzing the relationship between “information in the query”, “the inherent knowledge of the MLLM”, and “the knowledge retrieved” to allow “the knowledge retrieved” to correct the MLLM’s responses, fostering a reflective process. We have also attempted to use a single prompt to have the MLLM generate and then reflect on its answer, but it still gets influenced by the retrieved information.
In this way, we can fully utilize multimodal information and handle the fine-grained correspondences between them, enhancing the MLLMs’ ability to reason and answer questions in complex scenarios.
6. Experiment
6.1. Dataset
To validate the effectiveness of our method, we selected four representative KB-VQA datasets, each with its own focus areas:
  • OVEN (Hu et al., 2023b): An Open-domain Visual Entity Recognition dataset, primarily examining the ability to recognize the names of visual entities.
  • INFO SEEK (Chen et al., 2023): An extension of the OVEN dataset, focusing on the coarse-grained knowledge behind entities, environments, etc., in images. It requires identifying the image and then discovering the knowledge behind it.
  • OK-VQA (Marino et al., 2019): A classic KB-VQA dataset focusing on open-domain knowledge, featuring images paired with open-ended questions.
  • E-VQA (Yang et al., 2023): An event-centric dataset, primarily evaluating the ability to recognize events and the knowledge behind them.
Table 1 shows some characteristics of each dataset. The more stars in question granularity, the finer the question. The higher the popularity of knowledge, the more general it is, meaning the MLLM is more likely to have learned it during pre-training. Note that since our method is conducted in a zero-shot setting, we only selected the test sets of these datasets. Due to the large size of the original test sets for OVEN and INFO SEEK, we sampled some examples using an arithmetic sequence for testing, and the term ‘$s$’ is used in Table  1 and subsequent experimental results to indicate this.
Table 1. Characteristics of Different Dataset
Dataset Tests Number Knowledge Source Knowledge Granularity Knowledge Popularity
[…]
Table 2. Main results of the experiment. And indicates that the result is from experiments conducted on the full version of the test set, sourced respectively from Hu et al. (2023b) and Chen et al. (2023). Except for the SOTAs, which are trained ([Uncaptioned image]), other methods are performed in a zero-shot scenario ([Uncaptioned image]).
Model Dataset
OVENs INFO SEEKs OK-VQA E-VQA
SOTA[Uncaptioned image] 21.70 22.10 66.10 19.42
LlaVa NEXT-7b[Uncaptioned image] 9.51 6.37 73.33 10.51
LlaVa NEXT-7b + KU-RAG[Uncaptioned image] 10.80 9.09 73.07 11.00
Qwen 2.5-VL-32b[Uncaptioned image] 28.44 27.01 73.04 14.82
Qwen 2.5-VL-32b + KU-RAG[Uncaptioned image] 30.42 27.03 73.46 21.10
GPT-4o[Uncaptioned image] 22.30 36.05 75.52 15.17
GPT-4o + KU-RAG[Uncaptioned image] 26.50 38.35 77.23 26.16
6.2. Baseline
For the selection of baselines, we chose representative MLLMs with different parameter sizes. Due to variations in the formats and objectives of each dataset, there is no single unified state-of-the-art (SOTA) model across all of them. To ensure a fair comparison, we select the best-performing model for each dataset as its respective SOTA baseline.
  • SOTA: For OVEN dataset, we use PaLI-17B (Chen et al., 2022), as reported by Hu et al. (2023b). For INFO SEEK and OK-VQA datasets, the SOTA model is PaLI-X (Chen et al., 2022), as reported in the work of Chen et al. (2023). For E-VQA dataset, we adopt the best results of the SOTA model MuKEA (Ding et al., 2022), as reported Yang et al. (2023).
For the MLLMs:
  • GPT-4o: A closed-source MLLM launched by OpenAI with powerful multimodal content understanding and reasoning capabilities.
  • LlaVa NEXT-7b (Liu et al., 2024): The 7b parameter version of the latest LlaVa model, an open-source MLLM.
  • Qwen 2.5-VL-32b: A 32b-parameter open-source MLLM developed by Alibaba, capable of advanced vision-language understanding and reasoning.
6.3. Experimental Setting
Our experiments were conducted on RTX 4090 GPUs. GPT-4o and Qwen 2.5-VL-32b use the base version of the API interface, while LlaVa conducts experiments using the Hugging Face transformers library  [^1] https://huggingface.co/docs/transformers/main. The LLaVa NEXT-7b model used the weight file ‘llava-v1.6-mistral-7b-hf’. In our method’s settings, OVEN, INFO SEEK, and OK-VQA all use entities as the knowledge unit, while E-VQA uses events as the knowledge unit. For the recall of knowledge units and chunks, the top-k is set to 3. For the experiment evaluation, we used accuracy as the metric.
6.4. Main Result
As shown in Table  2, we have the following findings.
Zero-shot Capability of MLLMs. MLLMs exhibit remarkable zero-shot capabilities in image understanding and reasoning. Compared to the previous SOTA model for KB-VQA, GPT-4o shows improvements of 0.6%, 13.95%, and 9.42% on the OVEN, INFO SEEK, and OK-VQA datasets, respectively, benefiting from its extensive world knowledge acquired during pre-training. However, as the E-VQA dataset involves less commonly known news-related knowledge, GPT-4o struggles to outperform the trained SOTA model in this scenario.
Smaller models, such as LLaVA NEXT-7b and Qwen 2.5-VL-32b, also demonstrate strong zero-shot reasoning ability but fall short of GPT-4o. Notably, Qwen 2.5-VL-32b performs significantly better than LLaVA NEXT-7b across all datasets, highlighting its stronger multi-modal understanding.
Superior Performance of MLLM+KU-RAG. Our proposed method, MLLM+KU-RAG, achieves superior results across all datasets. In a zero-shot scenario, even without prior exposure to the training set knowledge, GPT-4o+KU-RAG outperforms the existing SOTA models by 4.8%, 16.25%, 11.13%, and 6.74% on the four datasets, respectively, validating its strong reasoning and retrieval augmentation capabilities.
Furthermore, Qwen 2.5-VL-32b+KU-RAG achieves the highest zero-shot accuracy among mid-sized models, surpassing LLaVA NEXT-7b+KU-RAG by a large margin. This suggests that models with better intrinsic multi-modal alignment can benefit more from knowledge retrieval.
Enhancement of MLLM by KU-RAG. Integrating KU-RAG with GPT-4o leads to performance gains of 4.2%, 2.3%, 1.7%, and 10.99% across the respective datasets. The most significant improvement is observed on E-VQA, where KU-RAG provides crucial missing knowledge, while the gain on OK-VQA is smaller due to the model’s inherent knowledge coverage.
For smaller models, the benefits of KU-RAG are also evident, though less pronounced than in GPT-4o. Qwen 2.5-VL-32b sees improvements of 1.98%, 2.1%, 0.4%, and 6.28% across the four datasets, showing its ability to integrate retrieved information effectively. In contrast, LLaVA NEXT-7b benefits the least from RAG, suggesting that both model architecture and parameter scale play key roles in utilizing external knowledge efficiently.
These results highlight that KU-RAG is particularly effective when paired with strong multi-modal models, significantly enhancing factual consistency and knowledge coverage in zero-shot settings.
Figure 5. The Results of Ablation Study
6.5. Ablation Study
To validate the effectiveness of each component in our proposed method, we designed ablation experiments comparing the following models:
  • $w/o$ KCC: This model omits the knowledge correction chain (KCC), relying instead on the model’s analysis of the question and the retrieved information in a single-turn Q&A setup.
  • $w/o$ KU: This model removes the fine-grained retrieval approach (i.e., knowledge unit), converting the information from images into captions and using a text-only retrieval modality.
Additionally, we included the full implementation of the GPT-4o+KU-RAG method, as well as a standalone GPT-4o. The experimental results are shown in Figure 5. From the figure, we can draw the following conclusions:
Effectiveness of GPT-4o+KU-RAG. The GPT-4o+KU-RAG method consistently achieves the highest performance across all four benchmarks, indicating the effectiveness and complementarity of its components. This result highlights the importance of integrating both KU and KCC in the retrieval-augmented generation pipeline.
Impact of Removing KCC. Removing KCC and using single-turn dialogue markedly reduces model performance across four datasets, with decreases of 6%, 18.79%, 8.3%, and 7.97%, respectively. Except for the E-VQA dataset, the model’s performance is inferior to using only GPT-4o. This likely occurs because the model struggles to effectively focus on the original question’s image and manage the logical relationships between the query information, its own knowledge, and the retrieved knowledge. Consequently, some questions that the model could originally answer correctly are answered incorrectly due to interference from the injected information.
Impact of Removing KU. Removing KU and adopting a coarse-grained, single-modality retrieval approach results in a slight performance drop across datasets, with the most significant decrease observed in the INFO SEEK dataset (9.04%). This is partly because INFO SEEK requires matching detailed image content and background knowledge, and converting the original image to captions loses a substantial amount of visual information. As illustrated by examples in Figure  1, it’s challenging to accurately match “Karnin Lift Bridge” using just the text “bridge," let alone find corresponding background knowledge. Furthermore, introducing incorrect knowledge adds noise, impeding the MLLM’s reasoning process and leading to erroneous results. The smallest performance drop is observed in the E-VQA dataset (1.65%), likely because, in this dataset, the images primarily serve to supplement information, allowing text-only retrieval to still achieve reasonably good matches.
Comparison with GPT-4o Only. Notably, the performance of certain ablation variants (particularly the version excluding KCC) underperforms the standalone GPT-4o model on datasets such as INFO SEEK and OK-VQA. This observation reinforces the critical importance of knowledge integration mechanisms: when retrieved content lacks proper orchestration through strategies like KCC or KU, the model may assimilate irrelevant or contradictory information that disrupts its reasoning processes. Our findings demonstrate that effective retrieval-augmented generation depends not only on external knowledge access but more fundamentally on systematic integration frameworks. Specifically, successful implementations require architectures that enable contextual coherence and logical synthesis of retrieved information during the reasoning phase, highlighting the necessity of deliberate fusion strategies over mere knowledge injection.
7. Related Work
7.1. Knowledge-based Visual Question Answering
Knowledge-based Visual Question Answering (KB-VQA) aims to leverage external knowledge to assist in answering questions about images  (Marino et al., 2019; Gardères et al., 2020). In early KB-VQA approaches (Zhu et al., 2020; Gao et al., 2022; Lin and Byrne, 2022), Wikipedia was often used as the external knowledge source for KB-VQA (Caffagni et al., 2024), leading to the common adoption of a retriever-reader framework. This framework first retrieves textual knowledge relevant to the question and image, and then “reads” the text to predict the answer. However, this passage dense retrieval method is a unimodal, coarse-grained text-to-text retrieval process, which struggles with specific, fine-grained questions such as visual entities (Hu et al., 2023b), events (Yang et al., 2023), and visual information-seeking questions (Chen et al., 2023).
Since the emergence of LLMs, some methods have explored using implicit knowledge from LLMs in addition to retrieving information from databases like Wikipedia (Caffagni et al., 2024). Typically, they convert images into tags or captions and then use GPT to retrieve related knowledge (Gui et al., 2021). However, there is a gap between the query and the LLM’s knowledge source. To address this,  Hu et al. (2023a) proposed a prompt-guided image captioning method that controls the visual entities in generated captions based on textual queries, replacing general captions with question-dependent ones. Although some methods attempt to mitigate the loss of visual information by incorporating visual features (Salaberria et al., 2023) or enriching prompts with candidate answers (Shao et al., 2023), they have not completely solved the issue of information loss. Recently, Hao et al. (2024) introduce a self-bootstrapped visual-language model that refines retrieved knowledge using a selector-answerer framework, significantly improving knowledge selection and QA accuracy, but their method still requires complex training.
Our approach shifts the focus from retrieval optimization to knowledge representation by introducing Knowledge Units, which serve as structured bridges between queries and multimodal knowledge. Rather than treating retrieval as an isolated process, we integrate multimodal information into a unified retrieval and reasoning framework.
7.2. Multimodal Retrieval-augmented Generation
Although LLMs possess strong general knowledge answering capabilities, they still face limitations when dealing with domain-specific knowledge, outdated information, and avoiding hallucinations  (Gao et al., 2023). To address these issues, Retrieval-Augmented Generation (RAG) was developed. RAG enhances the answering ability of LLMs by retrieving relevant document fragments from external knowledge bases. Specifically, the RAG approach involves multiple modules such as data storage, query optimization, document retrieval, and answer generation. The basic process matches user queries with documents from a large external knowledge base, retrieves relevant document fragments, and generates answers by integrating this information through a generation model. This process is similar to the knowledge retrieval mechanisms used in KB-VQA.
Building on this, RAG has evolved to optimize the retrieval and generation processes. For example, GRAG (Graph Retrieval-Augmented Generation) improves the relevance of information and generation quality by emphasizing the importance of subgraph  (Hu et al., 2024). The FiD-RAG (Fusion-in-Decoder RAG) model parallelly fuses multiple retrieved documents during the generation stage, allowing the model to comprehensively integrate background knowledge from different sources  (Izacard and Grave, 2020). Moreover, DPR-RAG (Dense Passage Retrieval RAG) introduces dense retrieval techniques that significantly improve retrieval accuracy, quickly locating highly relevant fragments from large document collections  (Karpukhin et al., 2020).
Unlike these approaches, which primarily refine retrieval mechanisms, our method focuses on the structured representation of multimodal knowledge. By organizing knowledge into Knowledge Units, we establish a persistent, query-aware knowledge structure, ensuring fine-grained, contextually relevant retrieval.
8. Conclusion
In this paper, we introduce the KU-RAG, aimed at enhancing MLLMs by incorporating fine-grained retrieval of domain-specific knowledge. To improve the effectiveness of retrieval, we propose the concept of “knowledge units”, which allows for more targeted access to relevant information. Furthermore, we design a knowledge correction chain strategy to verify and refine the retrieved knowledge, which can mitigate errors and hallucinations, enhancing the overall reliability and coherence of the generated answers in VQA tasks. Our experimental results demonstrate significant performance gains across multiple KB-VQA benchmarks, highlighting the effectiveness of our approach.
Future research directions may explore dynamic knowledge updates to improve adaptability, and integrate user feedback to enhance retrieval relevance and answer accuracy. Extending KU-RAG to other multimodal tasks like visual dialogue could further demonstrate its generalizability.